# Universal Solvability for Robot Motion Planning on Graphs

### v1

@@ Abstract
Finally, we consider the Graph Edge Augmentation for Universal Solvability (EAUS) problem where given a connected graph $G$ that is not universally solvable for $p$ robots, the question is to check if for a given budget $\beta$, at most $\beta$ edges can be added to $G$ to make it universally solvable for $p$ robots. We provide an upper bound of $p-2$ on $\beta$ for general graphs. On the other hand, we also provide examples of graphs that require $\Theta(p)$ edges to be added in order to achieve universal solvability for $p$ robots. Next, we consider the augmentation problem Graph Vertex and Edge Augmentation for Universal Solvability (VEAUS) involving the addition of $\alpha$ vertices along with the addition of $\beta$ many edges to make a graph universally solvable for $p$ robots, and we provide further lower bounds for $\alpha$ and $\beta$.
@@ 1 Introduction
The motivation for our work also stems from prior studies that investigate coordinated motion planning in restricted environments, such as trees [1, 18] and bi-connected graphs [16, 21]. These works provide efficient algorithms for checking the feasibility of configuration transitions on a given pair of $S$ and $T$; however, they typically do not address whether all configurations are mutually reachable. While group-theoretic approaches have been explored in [16] to examine graph classes for which any pair of configurations over the same vertex set are mutually reachable, the notion of universal solvability, a property that guarantees the system can adapt to any dynamic task assignment (i.e., can reconfigure from any configuration to any other), remains underexplored, especially in general graphs. In this paper, we aim to bridge this gap by analysing the USolR problem from both a structural and algorithmic perspective for general graphs.
@@ 1.1 Related Work
Robot motion planning and path finding have received significant attention, both in theoretical and applied contexts over the past few decades [13, 14, 19]. Closely related problems are: token swapping [3, 22] and pebble motion on graphs [9, 11, 15, 16, 20]. In the token swapping problem, the tokens are swapped on the endpoints of an edge whereas in the pebble motion problem, the pebbles move to a neighbouring unoccupied vertex. More recently, coordinated motion planning has been studied with respect to minimising various target objectives, such as the makespan of the schedule of the robots [2, 7, 23] and the total distance travelled by all robots [5, 7, 10, 23]. Problems related to robot motion planning arise in robotics [4], motion planing [6], puzzle design [17], and distributed systems [8], and have motivated extensive research on their complexity, solvability, and algorithmic strategies. Now, we discuss a few key contributions related to the feasibility variants of robot motion planning.
@@
Kornhauser, Miller, and Spirakis [16] study the Pebble Motion problem on graphs, where $p<n$ labelled pebbles are placed on distinct vertices of an $n$-vertex graph $G$, and the goal is to reach a target configuration via moves to adjacent unoccupied vertices. They reduce the problem to a permutation group setting, where valid move sequences form a group $R(P)$. For various graph classes (e.g., biconnected, separable, trees), $R(P)$ is shown to be transitive or decomposable into transitive components. The authors prove that $R(P)$ contains either the alternating group $\mathbf{A}_{p}$ or the symmetric group $\mathbf{S}_{p}$, depending on the structure of $G$, and show that when a generator of prime length is available, the diameter of the group is sub-exponential, linking algebraic properties with motion planning. Auletta, Monti, Parente, and Persiano [1] address the Pebble Motion problem on trees and present a linear-time algorithm that reduces it to the Pebble Permutation problem, where the set of occupied vertices remains the same and the goal is to realise a given permutation of the pebbles in these vertices through a sequence of valid moves. Their approach yields an $\mathcal{O}(n)$-time decision algorithm and an $\mathcal{O}(n+\ell)$-time constructive algorithm, where $\ell$ is the length of the move sequence and $n$ is the number of vertices.
@@
Yu and Rus [24] generalise the classical Pebble Motion problem to pebble motion with rotations (PMR) for $p$ pebbles on an $n$-vertex graph, where in addition to simple moves into unoccupied vertices ($p<n$), synchronous cyclic rotations along disjoint cycles are also allowed. In the fully packed case ($p=n$), only rotations are feasible. They model reachable configurations as elements of a subgroup $\mathbf{G}\leq\mathbf{S}_{n}$, generated by these rotations. For 2-edge connected but not 2-connected graphs, they prove that $\mathbf{G}$ contains either the alternating group $\mathbf{A}_{n}$ or the symmetric group $\mathbf{S}_{n}$, and establish an upper bound of $\text{diam}(\mathbf{G})=\mathcal{O}(n^{2})$, ensuring reachability in polynomial time. These results extend to general 2-edge-connected graphs and enable a unified framework where feasibility can be tested in linear time, and a complete plan can be computed in $\mathcal{O}(n^{3})$ time for all values of $p\leq n$.
@@
Goraly and Hassin [11] study the Multi-colour Pebble Motion problem, where pebbles of different colours must be rearranged to match a target colour configuration. They extend earlier work by analysing the coloured setting and provide a linear-time feasibility test. Their method reduces the problem to a tree with transshipment vertices (vertices that facilitate movement, however, cannot hold pebbles). Adapting the algorithm by Auletta et al. [1], they use equivalence classes of vertices to efficiently handle colour constraints and movement rules, resolving an open problem from earlier literature.
@@
Masehian and Nejad [18] investigate Multi-Robot Motion Planning on trees, aiming to determine whether any initial configuration can reach any target configuration without collisions. They introduce the notion of solvable trees, partially solvable trees, and minimal solvable trees, and analyse conditions under which global and local interchanges are feasible. Their approach uses the concept of influence zones and proposes a linear-time algorithm based on computing each robot’s maximum reachability space.
@@ 1.2 Our Contributions
The group theoretic results in the works of Yu and Rus [24] motivate us to look into equivalence classes of configurations that are reachable from each other. We prove that the sizes of all such equivalence classes are equal. This results in the fact that the number of reachable configurations is a factor of the number of all possible configurations when $p=|V|$. For $p\leq|V|$, we use a similar argument to arrive at the same conclusion. This result is used to design a simple and efficient randomized algorithm solving USolR.
@@
To analyse this problem structurally, we introduce a deterministic accumulation procedure in Section 3 that transforms any configuration to a canonical one where robots occupy a fixed vertex set $V_{p}\subseteq V$, $|V_{p}|=p$. This accumulation procedure is not used as a subroutine in the algorithm design but used for analyzing the algorithms later on. Recall that our definition of the equivalence relation over configurations is based on mutual reachability and we show that each equivalence class has the same size. Using this, we prove that in any instance that is not universally solvable, at least half of the configurations are unreachable from a fixed configuration.
@@
With this insight, we design a randomized algorithm that samples a configuration uniformly at random and uses a linear-time subroutine to test reachability from the identity configuration; this algorithm has a one-sided error and runs in $\mathcal{O}(|V|+|E|)$ time (Section 4). While the randomized algorithm is not directly used in the final deterministic solution, it is independently interesting due to its linear-time performance on general graphs. This is in sharp contrast to the best known deterministic algorithm for universal solvability, which requires quadratic time even for planar graphs [24].
@@
In Section 5, we derandomize the previous algorithm that reduces the number of required reachability checks from $p!$ (i.e., a trivial derandomization of the previous algorithm) to just $p-1$ by observing that reachability of adjacent transpositions (swaps between robots $t$ and $t+1$ for some $t$) suffices to check universal solvability.
@@
We also propose an optimized algorithm that takes into account structural graph properties (such as connectivity and the existence of sufficiently large 2-connected components). This algorithm runs in $\mathcal{O}(p\cdot(|V|+|E|))$ time in the worst case and improves to $\mathcal{O}(|E|)$ when the graph is sufficiently dense with $|E|=\omega(p\cdot|V|)$ (Section 6).
@@
Further, we consider the Graph Edge Augmentation for Universal Solvability (EAUS) problem where given a connected graph $G$ that is not universally solvable for $p$ robots, the question is to check if for a given budget $\beta$ at most $\beta$ edges can be added to $G$ to make it universally solvable for $p$ robots (Section 7). For a graph that is 2-edge-connected but not 2-connected, it may not necessarily be universally solvable for $p=|V|$ robots. However, we show that the addition of a single edge ensures universal solvability in such case. In case of simple cycles, we show that the addition of a chord suffices to render the given instance universally solvable for all $p\leq|V|$. In contrast, for 1-edge-connected graphs, the augmentation problem becomes more challenging. We prove that $\Theta(p)$ edges are required in the worst case for star graphs. On the positive side, we show that $\beta\leq p-2$ for all graphs.
@@
Finally, we consider the problem Graph Vertex and Edge Augmentation for Universal Solvability (VEAUS) where along with $\beta$ edge additions, we allow the addition of an $\alpha$ vertices to the input graph and ask if it is sufficient to make the graph universally solvable for $p$ robots (also in Section 7). We provide lower bounds for $\alpha$ and $\beta$ for this problem. An interesting consequence of these lower bounds is that there is an infinite class of graphs for which a budget of $\alpha=\Omega(\sqrt{|V|})$ is required, even when $\beta=\Theta(\sqrt{|V|})$, for universal solvability of a graph $G(V,E)$ with $p=|V|$ robots.
@@ 2 Preliminaries
We consider an underlying undirected graph $G(V,E)$ with $|V|=n$ and $|E|=m$. Consider a set $R$ of $p\in\mathbb{N}$ robots. For convenience, we assume the robots are named $1$, $2$, $\ldots$, $p$; hence $R=[p]$. We use both $R$ and $[p]$ to denote the set of robots, depending on the context. A configuration, $S:R\to V$ is an injective map from the set of robots to the set of vertices of the underlying graph $G(V,E)$; for robot $i$, $S(i)$ denotes the vertex occupied by it. The configuration is an injective map as it models the constraint that no two robots can occupy the same vertex in a configuration, i.e., $i\neq j\implies S(i)\neq S(j)$. We denote by $\mathcal{S}_{G}=\{S:R\to V\mid S\text{ is a configuration}\}$ the set of all configurations of the robots in $R$ that are possible on the graph $G$. Note that $|\mathcal{S}_{G}|=\frac{n!}{(n-p)!}$.
@@
  • Simple Path Move: A simple path move is defined as a pair of configurations that can be achieved from one another by sliding robots along a path synchronously, as shown in Figure 1. Formally, for configurations $S,S^{\prime}$, the pair $(S,S^{\prime})$ defines a simple path move if there is a path $P=(u_{0},u_{1},u_{2},\ldots,u_{k})$ in $G$ such that,
@@
  • Simple Rotation Move: A simple rotation move is a move along a simple cycle (all whose vertices are occupied by robots) such that all the vertices in that cycle synchronously move to their immediate neighbour in either clockwise or anticlockwise direction in one such move, as depicted in Figure 2. Formally, for configurations $S,S^{\prime}$, the pair $(S,S^{\prime})$ defines a simple rotation if there is a cycle $C=(u_{0},u_{1},u_{2},\ldots,u_{k}=u_{0})$ in $G$ such that,
@@
Further, for a configuration $S$ and a permutation $\pi:[p]\rightarrow[p]$, their composition $S\circ\pi$, corresponds to another configuration where robots are renamed according to the permutation $\pi^{-1}$; if robot $i$ occupies vertex $S(i)$ in $S$, then robot $\pi^{-1}(i)$ also occupies vertex $S\circ\pi(\pi^{-1}(i))=S(i)$ in the configuration $S\circ\pi$, as illustrated in Figure 3. We state an observation regarding valid moves.
@@ 2.1 Equivalence Relation based on Reachable Configurations
Since an equivalence relation provides a partition of the underlying set into disjoint equivalence classes, the reachability relation $\tilde{R}$ also partitions the set $\mathcal{S}_{G}$ into disjoint equivalence classes based on mutual reachability. Further, we can generalise Section 2 to the notion of reachability as follows, by the same argument of renaming of the robots.
@@
For configurations $S$ and $S^{\prime}$, $S\sim S^{\prime}$ if and only if $S\circ\pi\sim S^{\prime}\circ\pi$ for any permutation $\pi:[p]\to[p]$.
@@
Suppose $S\sim S^{\prime}$. Then there exists a sequence of configurations $(S_{0}=S,S_{1},\ldots,S_{t}=S^{\prime})$ such that $(S_{k-1},S_{k})$ is a valid move for all $k\in[t]$. From Section 2, it follows that the sequence $(S_{0}\circ\pi=S\circ\pi,S_{1}\circ\pi,\ldots,S_{t}\circ\pi=S^{\prime}\circ\pi)$ is such that $(S_{k-1}\circ\pi,S_{k}\circ\pi)$ is a valid move for all $k\in[t]$. Thus, $S\circ\pi\sim S^{\prime}\circ\pi$.

### v2

@@ Abstract
Finally, we consider the Graph Edge Augmentation for Universal Solvability (EAUS) problem where given a connected graph $G$ that is not universally solvable for $p$ robots, the question is to check if for a given budget $\beta$, at most $\beta$ edges can be added to $G$ to make it universally solvable for $p$ robots. We provide an upper bound of p2p-2 on $\beta$ for general graphs. On the other hand, we also provide examples of graphs that require $\Theta(p)$ edges to be added in order to achieve universal solvability for $p$ robots. Next, we consider the augmentation problem Graph Vertex and Edge Augmentation for Universal Solvability (VEAUS) involving the addition of $\alpha$ vertices along with the addition of $\beta$ many edges to make a graph universally solvable for $p$ robots, and we provide further lower bounds for $\alpha$ and $\beta$.
@@ 1 Introduction
The motivation for our work also stems from prior studies that investigate coordinated motion planning in restricted environments, such as trees [auletta1999linear, masehian2009solvability] and bi-connected graphs [kornhauser1984coordinating, wilson1974graph]. These works provide efficient algorithms for checking the feasibility of configuration transitions on a given pair of $S$ and $T$; however, they typically do not address whether all configurations are mutually reachable. While group-theoretic approaches have been explored in [kornhauser1984coordinating] to examine graph classes for which any pair of configurations over the same vertex set are mutually reachable, the notion of universal solvability, a property that guarantees the system can adapt to any dynamic task assignment (i.e., can reconfigure from any configuration to any other), remains underexplored, especially in general graphs. In this paper, we aim to bridge this gap by analysing the USolR problem from both a structural and algorithmic perspective for general graphs.
@@ 1.1 Related Work
Robot motion planning and path finding have received significant attention, both in theoretical and applied contexts over the past few decades [ismail2018survey, jahanshahi2018robot, qin2023review]. Closely related problems are: token swapping [bonnet2018complexity, yamanaka2015swapping] and pebble motion on graphs [fujita2015pebble, goraly2010multi, kato2021pebble, kornhauser1984coordinating, surynek2009application]. In the token swapping problem, the tokens are swapped on the endpoints of an edge whereas in the pebble motion problem, the pebbles move to a neighbouring unoccupied vertex. More recently, coordinated motion planning has been studied with respect to minimising various target objectives, such as the makespan of the schedule of the robots [7949109, eiben_et_al:LIPIcs.SoCG.2023.28, 7342901] and the total distance travelled by all robots [deligkas_et_al:LIPIcs.ICALP.2024.53, eiben_et_al:LIPIcs.SoCG.2023.28, 10.5555/3535850.3535905, 7342901]. Problems related to robot motion planning arise in robotics [chung2018survey], motion planing [demaine2019coordinated], puzzle design [loyd1959mathematical], and distributed systems [1612674], and have motivated extensive research on their complexity, solvability, and algorithmic strategies. Now, we discuss a few key contributions related to the feasibility variants of robot motion planning.
@@
Kornhauser, Miller, and Spirakis [kornhauser1984coordinating] study the Pebble Motion problem on graphs, where $p<n$ labelled pebbles are placed on distinct vertices of an $n$-vertex graph $G$, and the goal is to reach a target configuration via moves to adjacent unoccupied vertices. They reduce the problem to a permutation group setting, where valid move sequences form a group $R(P)$. For various graph classes (e.g., biconnected, separable, trees), $R(P)$ is shown to be transitive or decomposable into transitive components. The authors prove that $R(P)$ contains either the alternating group $\mathbf{A}_{p}$ or the symmetric group $\mathbf{S}_{p}$, depending on the structure of $G$, and show that when a generator of prime length is available, the diameter of the group is sub-exponential, linking algebraic properties with motion planning. Auletta, Monti, Parente, and Persiano [auletta1999linear] address the Pebble Motion problem on trees and present a linear-time algorithm that reduces it to the Pebble Permutation problem, where the set of occupied vertices remains the same and the goal is to realise a given permutation of the pebbles in these vertices through a sequence of valid moves. Their approach yields an $\mathcal{O}(n)$-time decision algorithm and an $\mathcal{O}(n+\ell)$-time constructive algorithm, where $\ell$ is the length of the move sequence and $n$ is the number of vertices.
@@
Yu and Rus [yu2014pebblemotiongraphsrotations] generalise the classical Pebble Motion problem to pebble motion with rotations (PMR) for $p$ pebbles on an $n$-vertex graph, where in addition to simple moves into unoccupied vertices ($p<n$), synchronous cyclic rotations along disjoint cycles are also allowed. In the fully packed case ($p=n$), only rotations are feasible. They model reachable configurations as elements of a subgroup $\mathbf{G}\leq\mathbf{S}_{n}$, generated by these rotations. For 2-edge connected but not 2-connected graphs, they prove that $\mathbf{G}$ contains either the alternating group $\mathbf{A}_{n}$ or the symmetric group $\mathbf{S}_{n}$, and establish an upper bound of $\text{diam}(\mathbf{G})=\mathcal{O}(n^{2})$, ensuring reachability in polynomial time. These results extend to general 2-edge-connected graphs and enable a unified framework where feasibility can be tested in linear time, and a complete plan can be computed in $\mathcal{O}(n^{3})$ time for all values of $p\leq n$.
@@
Goraly and Hassin [goraly2010multi] study the Multi-colour Pebble Motion problem, where pebbles of different colours must be rearranged to match a target colour configuration. They extend earlier work by analysing the coloured setting and provide a linear-time feasibility test. Their method reduces the problem to a tree with transshipment vertices (vertices that facilitate movement, however, cannot hold pebbles). Adapting the algorithm by Auletta et al. [auletta1999linear], they use equivalence classes of vertices to efficiently handle colour constraints and movement rules, resolving an open problem from earlier literature.
@@
Masehian and Nejad [masehian2009solvability] investigate Multi-Robot Motion Planning on trees, aiming to determine whether any initial configuration can reach any target configuration without collisions. They introduce the notion of solvable trees, partially solvable trees, and minimal solvable trees, and analyse conditions under which global and local interchanges are feasible. Their approach uses the concept of influence zones and proposes a linear-time algorithm based on computing each robot’s maximum reachability space.
@@ 1.2 Our Contributions
The group theoretic results in the works of Yu and Rus [yu2014pebblemotiongraphsrotations] motivate us to look into equivalence classes of configurations that are reachable from each other. We prove that the sizes of all such equivalence classes are equal. This results in the fact that the number of reachable configurations is a factor of the number of all possible configurations when $p=|V|$. For $p\leq|V|$, we use a similar argument to arrive at the same conclusion. This result is used to design a simple and efficient randomized algorithm solving USolR.
@@
To analyse this problem structurally, we introduce a deterministic accumulation procedure in Section˜3 that transforms any configuration to a canonical one where robots occupy a fixed vertex set $V_{p}\subseteq V$, $|V_{p}|=p$. This accumulation procedure is not used as a subroutine in the algorithm design but used for analyzing the algorithms later on. Recall that our definition of the equivalence relation over configurations is based on mutual reachability and we show that each equivalence class has the same size. Using this, we prove that in any instance that is not universally solvable, at least half of the configurations are unreachable from a fixed configuration.
@@
With this insight, we design a randomized algorithm that samples a configuration uniformly at random and uses a linear-time subroutine to test reachability from the identity configuration; this algorithm has a one-sided error and runs in $\mathcal{O}(|V|+|E|)$ time (Section˜4). While the randomized algorithm is not directly used in the final deterministic solution, it is independently interesting due to its linear-time performance on general graphs. This is in sharp contrast to the best known deterministic algorithm for universal solvability, which requires quadratic time even for planar graphs [yu2014pebblemotiongraphsrotations].
@@
In Section˜5, we derandomize the previous algorithm that reduces the number of required reachability checks from $p!$ (i.e., a trivial derandomization of the previous algorithm) to just p1p-1 by observing that reachability of adjacent transpositions (swaps between robots $t$ and $t+1$ for some $t$) suffices to check universal solvability.
@@
We also propose an optimized algorithm that takes into account structural graph properties (such as connectivity and the existence of sufficiently large 2-connected components). This algorithm runs in $\mathcal{O}(p\cdot(|V|+|E|))$ time in the worst case and improves to $\mathcal{O}(|E|)$ when the graph is sufficiently dense with $|E|=\omega(p\cdot|V|)$ (Section˜6).
@@
Further, we consider the Graph Edge Augmentation for Universal Solvability (EAUS) problem where given a connected graph $G$ that is not universally solvable for $p$ robots, the question is to check if for a given budget $\beta$ at most $\beta$ edges can be added to $G$ to make it universally solvable for $p$ robots (Section˜7). For a graph that is 2-edge-connected but not 2-connected, it may not necessarily be universally solvable for $p=|V|$ robots. However, we show that the addition of a single edge ensures universal solvability in such case. In case of simple cycles, we show that the addition of a chord suffices to render the given instance universally solvable for all $p\leq|V|$. In contrast, for 1-edge-connected graphs, the augmentation problem becomes more challenging. We prove that $\Theta(p)$ edges are required in the worst case for star graphs. On the positive side, we show that $\beta\leq p-2$ for all graphs.
@@
Finally, we consider the problem Graph Vertex and Edge Augmentation for Universal Solvability (VEAUS) where along with $\beta$ edge additions, we allow the addition of an $\alpha$ vertices to the input graph and ask if it is sufficient to make the graph universally solvable for $p$ robots (also in Section˜7). We provide lower bounds for $\alpha$ and $\beta$ for this problem. An interesting consequence of these lower bounds is that there is an infinite class of graphs for which a budget of $\alpha=\Omega(\sqrt{|V|})$ is required, even when $\beta=\Theta(\sqrt{|V|})$, for universal solvability of a graph $G(V,E)$ with $p=|V|$ robots.
@@ 2 Preliminaries
We consider an underlying undirected graph $G(V,E)$ with $|V|=n$ and $|E|=m$. Consider a set $R$ of $p\in\mathbb{N}$ robots. For convenience, we assume the robots are named $1$, 22, $\ldots$, $p$; hence $R=[p]$. We use both $R$ and $[p]$ to denote the set of robots, depending on the context. A configuration, $S:R\to V$ is an injective map from the set of robots to the set of vertices of the underlying graph $G(V,E)$; for robot $i$, $S(i)$ denotes the vertex occupied by it. The configuration is an injective map as it models the constraint that no two robots can occupy the same vertex in a configuration, i.e., $i\neq j\implies S(i)\neq S(j)$. We denote by $\mathcal{S}_{G}=\{S:R\to V\mid S\text{ is a configuration}\}$ the set of all configurations of the robots in $R$ that are possible on the graph $G$. Note that $|\mathcal{S}_{G}|=\frac{n!}{(n-p)!}$.
@@
  • Simple Path Move: A simple path move is defined as a pair of configurations that can be achieved from one another by sliding robots along a path synchronously, as shown in Figure˜1. Formally, for configurations $S,S^{\prime}$, the pair $(S,S^{\prime})$ defines a simple path move if there is a path $P=(u_{0},u_{1},u_{2},\ldots,u_{k})$ in $G$ such that,
@@
  • Simple Rotation Move: A simple rotation move is a move along a simple cycle (all whose vertices are occupied by robots) such that all the vertices in that cycle synchronously move to their immediate neighbour in either clockwise or anticlockwise direction in one such move, as depicted in Figure˜2. Formally, for configurations $S,S^{\prime}$, the pair $(S,S^{\prime})$ defines a simple rotation if there is a cycle $C=(u_{0},u_{1},u_{2},\ldots,u_{k}=u_{0})$ in $G$ such that,
@@
Further, for a configuration $S$ and a permutation $\pi:[p]\rightarrow[p]$, their composition $S\circ\pi$, corresponds to another configuration where robots are renamed according to the permutation $\pi^{-1}$; if robot $i$ occupies vertex $S(i)$ in $S$, then robot $\pi^{-1}(i)$ also occupies vertex $S\circ\pi(\pi^{-1}(i))=S(i)$ in the configuration $S\circ\pi$, as illustrated in Figure˜3. We state an observation regarding valid moves.
@@ 2.1 Equivalence Relation based on Reachable Configurations
Since an equivalence relation provides a partition of the underlying set into disjoint equivalence classes, the reachability relation $\tilde{R}$ also partitions the set $\mathcal{S}_{G}$ into disjoint equivalence classes based on mutual reachability. Further, we can generalise Section˜2 to the notion of reachability as follows, by the same argument of renaming of the robots.
@@
For configurations $S$ and $^{\prime}, SS$$\sim S^{\prime}$ if and only if $S\circ\pi\sim S^{\prime}\circ\pi$ for any permutation $\pi:[p]\to[p]$.
@@
Suppose $S\sim S^{\prime}$. Then there exists a sequence of configurations $(S_{0}=S,S_{1},\ldots,S_{t}=S^{\prime})$ such that $(S_{k-1},S_{k})$ is a valid move for all $k\in[t]$. From Section˜2, it follows that the sequence $(S_{0}\circ\pi=S\circ\pi,S_{1}\circ\pi,\ldots,S_{t}\circ\pi=S^{\prime}\circ\pi)$ is such that $(S_{k-1}\circ\pi,S_{k}\circ\pi)$ is a valid move for all $k\in[t]$. Thus, $S\circ\pi\sim S^{\prime}\circ\pi$.